average senior high annual costs of tuition fee for all private schools last year was php43,700,a random sample costs this year for 45 private schools indicated that the sample mean was php45,800 and a sample standard deviation was php5,600 at 0.01 level of significance is there sufficent evidence for conclude that the cost increased ?

1
Expert's answer
2022-04-04T14:28:34-0400

"\u03bc=43700"

n=45

X=45800

"\sigma=5600"

"Z=\frac{X-\mu}{\frac{\sigma}{\sqrt{n}}}=\frac{45800-43700}{\frac{5600}{\sqrt{45}}}=2.52"

"Z_{\lambda/2}=Z_{0.01/2}=Z_{0.005}=2.807"

2.52<2.807


There is no sufficient evidence to conclude that the cost has increased.